There are two unopened boxes of skittles and M&M's, each with equally the same amount of candy.

There are 6 different colors of M&M's, each with varying amounts, Blue(24%), Red(26%), Purple(23%), Green(13%), and Orange(6%), and Yellow(8%)

There are 6 different colors of Skittles as well, with there being an equal chance of getting each color(1/6)

What is the probability of getting a skittle?

What is P(Skittle l Blue)?
Answer #1

1)probability of getting a skittle =1/2 (as there are equally the same amount of candy. in both the boxes)

2)

P(Blue) =P(Skittle and blue)+P(M&M and Blue)=(1/2)*(1/6)+(1/2)*(0.24)=61/300

hence P(Skittle|Blue)= P(Skittle and blue)/P(blue) =(1/2)*(1/6)/(61/300)=25/61